# Embracing Large Language Models in Traffic Flow Forecasting

### v1

@@ 3.3 Large Language Model Based Selector
3.4 Large Language Model Enhanced Prediction
The LEAF framework consists of a predictor and a selector. The predictor (the graph branch and the hypergraph branch) is pre-trained on the training set. During test time, the predictor first predicts, and the selector then selects. The selection results are used to supervise the predictor, and thus the two modules benefit from each other, achieving large language model enhanced prediction. Concretely, given the input data $\bm{X}\in\mathbb{R}^{T\times N\times F}$, the predictor generates two forecasts, i.e. $\bm{Y}^{G}$ and $\bm{Y}^{H}$. Subsequently, the selector constructs choice sets and use the LLM to find the best option for each individual vertex. The selection results is denoted as $\hat{\bm{y}}_{i}$, $i\in\mathcal{V}$, which are then used to supervise the predictor. Conceivably, $\hat{\bm{y}}_{i}$ may not be the optimal choice in the choice set, and therefore, we adopt a ranking loss described as follows:
@@ 3.4 Large Language Model Enhanced Prediction
By supervising the predictor, the two modules (i.e. the predictor and the selector) benefit from each other. A better predictor yields better choices, which benefits the selector; better selection results provide better supervision signals for the predictor. Through the iteration of prediction and selection, we can achieve large language model enhanced prediction. The LEAF framework during inference is summarized in Algorithm 1.
@@ 4.1 Experimental Setup
Datasets. We adopt three widely used datasets in traffic flow forecasting, including PEMS03, PEMS04, and PEMS08. The datasets are publicly available and collected by California Transportation Agencies (CalTrans) Performance Measurement Systems (PEMS).  [^1] https://pems.dot.ca.gov/. The traffic data come from sensors on the road of various places in California, and they are counted every five minutes.
Evaluation Metrics. We follow standard setting Li et al. (2018) that use one-hour historical data (i.e. $T=12$ timesteps) to forecast one-hour future (i.e. $T^{\prime}=12$ timesteps). For evaluation the prediction results, we use three standard metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE) Song et al. (2020).
Baseline Methods. We compare LEAF with a variety of baseline methods, including DCRNN Li et al. (2018), ASTGCN Guo et al. (2019), STSGCN Song et al. (2020), HGCN Wang et al. (2022), DyHSL Zhao et al. (2023), STAEformer Liu et al. (2023), COOL Ju et al. (2024), and LLM-MPE Liang et al. (2024). Among these methods, ASTGCN, STSGCN, and COOL are based on graph neural networks, while HGCN and DyHSL are based on hypergraph neural networks. STAEformer directly takes the transformer architecture for the traffic data, while LLM-MPE use LLMs to understand and predict traffic data with textual guidance.
Implementation Details. For the dual-branch predictor, the number of layers $L$ for both branches is set to 7. We use linear mapping as the input embedding, and the hidden dimension $d$ is set to 64, which is also shared across all baselines. We also use a two layer MLP to map the last hidden embedding to the output. In choice set construction, smoothing is implemented with an average filter, upward/downward trend increases/decreases the traffic flow by 1% to 12% (12 timesteps, linearly increasing), overestimate/underestimate increases/decreases the traffic flow by 5% (for all 12 timesteps). As for the loss function in Eq. 7, we adopt Huber distance for $\Delta(\cdot,\cdot)$ and the margin $\epsilon$ is set to 0. When training with this loss function, we update the parameters for $M$ iterations, which is set to 5. For the prediction-selection loop, we set $K$ to 2. For the LLM, we use LLaMA 3 70B AI@Meta (2024) and vLLM Kwon et al. (2023) as the inference software.
@@ 4.2 Main Results
The performance of LEAF in comparison with baselines are shown in Table 1. According to the results, we have several observations:
Firstly, the proposed LEAF achieves a consistent lead in all three datasets, which demonstrates the effectiveness of the proposed framework that adopts a dual-branch traffic flow predictor and a LLM-based selector. Prior methods often fail to provide satisfactory predictions when there is test-time distribution changes, as they cannot learn adaptively, leaving the room for improvement.
Secondly, the method that utilizes the generative ability of LLMs (i.e. LLM-MPE) does not perform well on all datasets. As we can see on PEMS03 and PEMS04 datasets, its predictions are generally worse or similar compared to simple methods using graph neural networks. Since LLMs are not very adept in capturing complex spatio-temporal relations, it is reasonable that we see such results on datasets with larger networks like PEMS03.
@@ 4.3 Ablation Study
We also perform ablation studies on the PEMS08 dataset, and the results are shown in Table 2. Specifically, we perform the following experiments. E1 measures the performance of graph branch only without the LLM-based selector. E2 measures the performance of the hypergraph branch without the selector. The vanilla version (E1 and E2) of both branches performs much worse than LEAF. E3 uses the graph branch in conjunction with the selector, which leads to performance degradation compared to LEAF. This suggests that non-pair-wise relations are important in traffic flow forecasting. E4 only uses the hypergraph branch together with the selector. Similarly, it performs worse than LEAF, which shows that pair-wise and non-pair-wise relations are both important. Moreover, E3 and E4 perform better than E1 and E2, which shows the effect of the LLM-based selector. E5 removes the transformations, i.e. $\mathcal{T}=\emptyset$. This leads to slightly worse performance, showing the effectiveness of providing more choices. E6 removes the ranking loss, which means that the predictor are not further trained with the results from the selector (reducing to $K=1$). This experiment demonstrates the effectiveness of supervising the predictor with a ranking loss using the selection results.
@@ 4.4 Hyper-parameter Analysis
We perform experiments with respect to two hyper-parameters: (a) $M$, i.e. the number of iterations when training with the ranking loss in Eq.7 and Eq.8, and (b) the number of $K$, i.e. the number of prediction-selection iterations in Algorithm 1. The results on the PEMS08 dataset are shown in Figure 4. As can be seen from the figure, when $M$ increases, both MAE and RMSE decreases, and then plateaus after around 5. This shows that the predictor converges and therefore, we set $M$ to 5. For another hyper-parameter $K$, the optimal performance is achieved when $K$ is set to 2, and with more iterations, the error increases. One reason is that we use the same prompt across iterations without contexts of previous selections to save computation, so the same factor may be considered multiple times, leading to decreased accuracy. In practice, $K=2$ is good enough without too much computation for the LLM-based selector.
@@ 4.5 Visualization
Visualization of the Selection. We first provide visualizations of the historical traffic flow data, the choices in $\mathcal{C}_{i}$, and the option selected by the LLM-based selector (marked with a green tick). The results on the PEMS03 dataset is shown in Figure 5. Moreover, we also provide the analysis of the LLM-based selector, which is shown in Figure 6. According to the results, we can see that the traffic flow is generally going downward, since it is the end of the rush hours. In Figure 5, we can see that the LLM-based selector chooses the option with the lowest traffic flows. From its analysis in Figure 6, we can see that it understands that the time period to forecast is around the end of the rush hours, which is the reason why it selects the lowest option. This suggests that the LLM-based selector is able to understand changing traffic conditions.
Figure 7: The Mean Absolute Error (MAE) under different timesteps. LEAF better reduces long term errors.
Visualization of Errors in Different Timesteps. We then provide visualization of the mean absolute error (MAE) under different forecasting timesteps in Figure 7. The experiments are performed on the PEMS03 dataset, where we compare the MAE values of our framework compared to its two branches (i.e. the graph branch and the hypergraph branch). The results show that LEAF reduces forecasting errors generally. Although the errors are similar in the first few timesteps, LEAF quickly diverges from the two branches, resulting in significantly lower errors in the long term. This suggests that with the help of the discriminative ability of the LLM-based selector, our method can select predictions that are more accurate in the long run.
@@ 5 Conclusion
In this paper, we propose a novel framework named Large Language Model Enhanced Traffic Flow Predictor (LEAF), which consists of a dual-branch traffic flow predictor and an LLM-based selector. The predictor adopts two branches: the graph branch and the hypergraph branch, capturing the pair-wise and non-pair-wise spatio-temporal relations respectively. The selector uses the discriminative ability of the LLM to choose the best forecast of predictor. The selection results are then used to supervise the predictor with a ranking loss. We perform extensive experiments to demonstrate the effectiveness of the proposed LEAF framework.
@@
One limitation of this work is that we only focus on the traffic flow forecasting domain, due to the scope of this paper and the limited computational resources. In the future, we plan to extend this framework to more generalized spatio-temporal forecasting problem. Besides, we do not evaluate our methods on other traffic flow forecasting datasets due to limited resources. For future works, we plan to evaluate our framework using other datasets.

### v2

@@ 3.2 LLM-Based Selector
3.3 LLM-Enhanced Prediction
The LEAF framework consists of a predictor and a selector. The predictor (the graph branch and the hypergraph branch) is pre-trained on the training set. During test time, the predictor first predicts, and the selector then selects. The selection results are used to supervise the predictor, and thus the two modules benefit from each other, achieving LLM-enhanced prediction. Concretely, given the input data $\bm{X}\in\mathbb{R}^{T\times N\times F}$, the predictor generates two forecasts, i.e. $\bm{Y}^{G}$ and $\bm{Y}^{H}$. Subsequently, the selector constructs choice sets and uses the LLM to find the best option for each individual vertex. The selection results are denoted as $\hat{\bm{y}}_{i}$, $i\in\mathcal{V}$, which are then used to supervise the predictor. Conceivably, $\hat{\bm{y}}_{i}$ may not be the optimal choice in the choice set, and therefore, directly using MAE or MSE losses may lead to noises in the supervision signals. By comparison, the ranking loss Weinberger and Saul (2009); Sohn (2016) only requires that the positive candidate (the ones chosen by the LLM) is better than the negative candidates (the ones not chosen), described as follows:
@@ 3.3 LLM-Enhanced Prediction
By supervising the predictor, the two modules (i.e. the predictor and the selector) benefit from each other. A better predictor yields better choices, which benefits the selector; better selection results provide better supervision signals for the predictor. Through the iteration of prediction and selection, we can achieve LLM-enhanced prediction. The LEAF framework during inference is summarized in Algorithm 1.
@@ 4.1 Experimental Setup
Datasets. We adopt three widely used datasets in traffic flow forecasting, including PEMS03, PEMS04, and PEMS08. The datasets are publicly available and collected by California Transportation Agencies (CalTrans) Performance Measurement Systems (PEMS)  [^1] https://pems.dot.ca.gov/. These records come from sensors on the roads of various places in California, and they are counted every five minutes.
Evaluation Metrics. We follow standard setting Li et al. (2018) that use one-hour historical data (i.e. $T=12$ timesteps) to forecast one-hour future (i.e. $T^{\prime}=12$ timesteps). To signify the test-time distribution shift, we adopt small training sets of 10% of data and another 10% for validation. We choose a subset of non-overlapping slices in the test set. We adopt three standard metrics for evaluation, i.e., Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE) Song et al. (2020).
Baseline Methods. We compare LEAF with a variety of baselines, including DCRNN Li et al. (2018), ASTGCN Guo et al. (2019), STSGCN Song et al. (2020), HGCN Wang et al. (2022), DyHSL Zhao et al. (2023), STAEformer Liu et al. (2023), COOL Ju et al. (2024), and LLM-MPE Liang et al. (2024). Among these methods, DCRNN is a combination of GNN and RNN. ASTGCN, STSGCN, and COOL are based on GNNs, while HGCN and DyHSL are based on hypergraph neural networks. STAEformer uses the transformer architecture, while LLM-MPE uses LLMs to understand and predict traffic data with textual guidance.
Implementation Details. For the dual-branch predictor, the number of layers $L$ for both branches is set to 7. We use linear mapping as the input embedding, and the hidden dimension $d$ is set to 64, which is also shared across all baselines. We also use a two-layer MLP to map the last hidden embedding to the output. In choice set construction, smoothing is implemented with an average filter, upward/downward trend increases/decreases the traffic flow by 1% to 12% (12 timesteps, linearly increasing), overestimate/underestimate increases/decreases the traffic flow by 5% (for all 12 timesteps). As for the loss function in Eq. 7, we adopt Huber distance for $\Delta(\cdot,\cdot)$ and the margin $\epsilon$ is set to 0. When training with this loss function, we update the parameters for $M$ iterations, where $M$ is set to 5. For the prediction-selection loop, we set $K$ to 2. For the LLM, we use LLaMA 3 70B AI@Meta (2024) and vLLM Kwon et al. (2023) as the inference software.
@@ 4.2 Main Results
The performance of LEAF in comparison with prior methods is shown in Table 1. According to the metrics in the table, we have several observations:
Firstly, the proposed LEAF demonstrates consistent improvement in all three datasets, showing the effectiveness of the proposed framework that adopts a dual-branch traffic flow predictor and an LLM-based selector. Smaller models often fail to provide satisfactory predictions under test-time distribution changes, as they fall short in reason and generalization, leaving room for improvement.
Secondly, the methods that utilize the generative ability of LLMs (i.e. LLM-MPE) do not perform well on all datasets. As we can see on the PEMS03 and PEMS04 datasets, their predictions are generally worse or similar compared to simple methods using graph neural networks (e.g. STSGCN, COOL). Since LLMs are not adept at capturing complex spatio-temporal relations, it is reasonable that we see such results on datasets with larger networks like PEMS03.
@@ 4.3 Ablation Study
We also perform ablation studies on the PEMS08 dataset, and the results are shown in Table 2. Specifically, we perform the following experiments. E1 measures the performance of the graph branch only without the LLM-based selector. E2 measures the performance of the hypergraph branch without the selector. The vanilla version (E1 and E2) of both branches performs much worse than LEAF. E3 uses the graph branch in conjunction with the selector, which leads to performance degradation compared to LEAF. This suggests that non-pair-wise relations are important in traffic flow forecasting. E4 only uses the hypergraph branch together with the selector. Similarly, it performs worse than LEAF, which shows that pair-wise and non-pair-wise relations are both important. Moreover, E3 and E4 perform better than E1 and E2, which shows the effect of the LLM-based selector. E5 removes the transformations, i.e. $\mathcal{T}=\emptyset$. This leads to slightly worse performance, showing the effectiveness of providing more choices. E6 removes the ranking loss, which means that the predictor is not further trained with the results from the selector (reducing to $K=1$). This experiment demonstrates the effectiveness of supervising the predictor with a ranking loss using the selection results.
@@ 4.4 Hyper-parameter Analysis
We perform experiments with respect to two hyper-parameters: (a) $M$, i.e. the number of iterations when training with the ranking loss in Eq. 7 and Eq. 8, and (b) the number of $K$, i.e. the number of prediction-selection iterations in Algorithm 1. The results on the PEMS08 dataset are shown in Figure 4. As can be seen from the figure, when $M$ increases, both MAE and RMSE decrease, and then plateau after around 5. This shows that the predictor converges, and therefore, we set $M$ to 5. For another hyper-parameter $K$, the optimal performance is achieved when $K$ is set to 2, and with more iterations, the error increases. One reason is that we use the same prompt across iterations without contexts of previous selections to save computation, so the same factor may be considered multiple times, leading to decreased accuracy. In practice, $K=2$ is good enough without too much computation for the LLM-based selector.
@@
Figure 7: The Mean Absolute Error (MAE) under different timesteps. LEAF better reduces long-term errors.
Figure 8: The details about the prompt.
@@ 4.5 Visualization
Visualization of the Selection. We first provide visualizations of the historical traffic flow data, the choices in $\mathcal{C}_{i}$, and the option selected by the LLM-based selector (marked with a green tick). The results on the PEMS03 dataset are shown in Figure 5. Moreover, we also provide the analysis of the LLM-based selector, which is shown in Figure 6. According to the results, we can see that the traffic flow is generally going downward, since it is the end of the rush hour. In Figure 5, we can see that the LLM-based selector chooses the option with the lowest traffic flows. From its analysis in Figure 6, we can see that it understands that the time period to forecast is around the end of the rush hours, which is the reason why it selects the lowest option. This suggests that the LLM-based selector is able to understand changing traffic conditions.
Visualization of Errors in Different Timesteps. We then provide visualization of the mean absolute error (MAE) under different forecasting timesteps in Figure 7. The experiments are performed on the PEMS03 dataset, where we compare the MAE values of our framework to its two branches (i.e. the graph branch and the hypergraph branch). The results show that LEAF reduces forecasting errors generally. Although the errors are similar in the first few timesteps, LEAF quickly diverges from the two branches, resulting in significantly lower errors in the long term. This suggests that with the help of the discriminative ability of the LLM-based selector, our method can select predictions that are more accurate in the long run.
Visualization of the Prompt. We provide an example of the prompt in Figure 8. The prompt contains general information about the task and the traffic data, spatio-temporal information, the historical data, the task instructions, and the choice set constructed in Eq. 6. Additionally, we also describe the transformations applied to the prediction of the dual-branch predictor. These sources of information will be helpful for the LLM to decide the most appropriate choice.
@@ 5 Conclusion
In this paper, we propose a novel framework named Large Language Model Enhanced Traffic Flow Predictor (LEAF), consisting of a dual-branch traffic flow predictor and an LLM-based selector. The predictor adopts two branches: the graph branch and the hypergraph branch, capturing the pair-wise and non-pair-wise relations, respectively. The selector uses the discriminative ability of the LLM to choose the best forecast of the predictor. The selection results are then used to supervise the predictor. We perform extensive experiments to demonstrate the effectiveness of LEAF.
@@
One limitation of this work is that we only focus on the traffic flow forecasting domain, due to the scope of this paper and the limited computational resources. However, this framework can be extended to more generalized spatio-temporal forecasting problems. Besides, we have not evaluated our methods on other traffic flow forecasting datasets due to limited resources.
Another limitation is that the LLM is not fine-tuned during the process. The LLM selector in the LEAF framework can be further optimized throughout the process. Since the traffic datasets are relatively small, a potential solution is to use parameter-efficient fine-tuning strategies, including LoRA, adapter layers, or prefix-tuning to enhance the performance of the LLM selector. More specifically, given the final prediction results from the iterative prediction-selection cycles, the LLM is then fine-tuned with the answer (in the choices) nearest to the final prediction. This will potentially enhance the ability of the LLM selector.
Acknowledgment
This paper is partially supported by grants from the National Key Research and Development Program of China with Grant No. 2023YFC3341203 and the National Natural Science Foundation of China (NSFC Grant Number 62276002). The authors are grateful to the anonymous reviewers for critically reading this article and for giving important suggestions to improve this article.